find the four consecutive integers such that the sum of the second and the third is 68.
Answer by rapaljer(4671)

Let x = first
x+1 = second
x+2 = Third
x+3 = fourth

Second + Third = 68
x+1 + x+2 = 68
2x + 3 = 68
2x = 65
x= 32.5 Not an integer!

There is NO SOLUTION!!

For additional explanation on Word Problems, please see my own website. Look for the easy-to-remember and easy-to-spell link www.mathinlivingcolor.com. At the bottom of this page is a link that takes you to my Homepage. Near the top of the Homepage, look for the link: "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra" and look in Chapter 1, for Section 1.10 Word Problems.

You will probably like my explanations, especially the PRICE of them!! They are ALL FREE!! My own students before my retirement thought my books were easier to understand than the traditional books, and many of the solutions are given IN COLOR in the attached "Math in Living Color" pages.

Most any algebra topic you need is probably there somewhere. If you need help finding the right page on the website, please send me an Email at rapaljer@seminolestate.edu. If you like my website, please recommend it to family and friends.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida